# Replay Consolidation with Label Propagation for Continual Object Detection

### v1

@@ Abstract
Object Detection is a highly relevant computer vision problem with many applications such as robotics and autonomous driving. Continual Learning (CL) considers a setting where a model incrementally learns new information while retaining previously acquired knowledge. This is particularly challenging since Deep Learning models tend to catastrophically forget old knowledge while training on new data. In particular, Continual Learning for Object Detection (CLOD) poses additional difficulties compared to CL for Classification. In CLOD, images from previous tasks may contain unknown classes that could reappear labeled in future tasks. These missing annotations cause task interference issues for replay-based approaches. As a result, most works in the literature have focused on distillation-based approaches. However, these approaches are effective only when there is a strong overlap of classes across tasks. To address the issues of current methodologies, we propose a novel technique to solve CLOD called Replay Consolidation with Label Propagation for Object Detection (RCLPOD). Based on the replay method, our solution avoids task interference issues by enhancing the buffer memory samples. Our method is evaluated against existing techniques in CLOD literature, demonstrating its superior performance on established benchmarks like VOC and COCO.
Keywords Continual Learning  $\cdot$ Object Detection
Figure 1: Continual Learning for Object Detection pipeline. The model learns to detect new classes at each incremental training stage - corresponding to a new task. The additional issue is the missing annotations; the labeled example in the current task might have been learned as background in previous tasks (e.g., the class dog is labeled in Task 1 but not in Task 2 or 3.).
@@ 1 Introduction
Object Detection is a critical computer vision problem involving both the localization and classification of objects within an image, with applications from robotics to autonomous driving [1]. However, a key challenge of Deep Learning systems is that they cannot learn incrementally over time, as they are prone to Catastrophic Forgetting [2, 3] when training on new data. Continual Learning (CL) addresses this problem, introducing techniques to learn new information over time while retaining previously acquired knowledge.
Many CL studies focus on image classification, where replay-based approaches are the most common way to achieve strong results [4, 5, 6, 7]. These approaches involve storing a subset of samples from previous tasks in a memory buffer and replaying them when training on new data. However, Continual Learning for Object Detection (CLOD) is more complex because images seen in previous tasks may contain unlabeled classes that the model must learn as new classes in future tasks, causing missing annotations and training conflicts [8]. For example, the image in Fig. 1 has the classes person, dog, and car. However, this image has only one labeled class for each task. Then, this image could be saved multiple times in the replay memory, each time with different annotations based on the task, causing interference. Because of this, most of the CLOD literature is based on distillation-based approaches [8]. However, these approaches encounter several challenges, particularly the tendency to forget older classes when their instances are underrepresented in new training images.
To address the limitations of replay approaches for CLOD, we propose a novel technique called Replay Consolidation with Label Propagation for Object Detection (RCLPOD). Our approach consists of multiple components, including a technique called Label Propagation [9], a selection mechanism called Optimizing Class Distribution in Memory (OCDM) [10], a Feature Distillation technique [11], and a Masking Loss [12]. The key idea of our approach is to significantly reduce the task interference issue of the replay memory via the Label Propagation technique, enabling more efficient use of the samples saved in the replay memory. The superior performance of our approach is validated on the well-known CLOD benchmarks based on VOC and COCO [8].
While in the literature some CL studies were conducted either on YOLOv3 [12] and v5 [13], we test our approach on YOLOv8 [14], a recent and anchor-free version of the real-time object detector YOLO [15]. Therefore, this work can serve as a valuable benchmark for future research on contemporary YOLO architectures within the CLOD framework.
@@
  • We propose a novel approach called RCLPOD, which addresses the main issues of replay techniques in the CLOD setting.
  • We conduct extensive experiments on RCLPOD, demonstrating its superior performance on the established VOC and COCO CLOD benchmarks.
  • We benchmark our approach and other CL techniques using more modern versions of YOLO compared to other works in the literature.
The paper is structured as follows. Sec. 2 provides an overview of the CL literature focusing on CLOD. In Sec. 3, we describe our novel approach, RCLPOD. Sec. 4 describes the experimental setting, including the considered metrics, evaluated scenarios, and experimental details. In Sec. 5, we show and discuss the results. Finally, Sec. 6 concludes this work by discussing limitations and future research directions.
@@ 2 Related Work
Continual Learning aims to enable DL models to continuously accumulate knowledge over time without the need to re-train from scratch and with limited access to past data [16]. In particular, this framework has emerged to tackle the Catastrophic Forgetting problem [2, 3]; after a DL model is trained on one task and subsequently trained on another, it tends to forget the previously learned task. The methods in the CL literature can be grouped into three big families: rehearsal, regularization, and architecture-based approaches.
Most of the literature on Continual Learning for Object Detection focuses on the Class Incremental Learning (CIL) scenario, where each new task introduces new classes [8]. For example, Fig. 1 depicts a CLOD setting composed of three tasks where a new class must be learned each time. At the end of the training, the model should be able to provide predictions on all these classes. Moreover, most of the works were conducted on two-stage object detectors, whose time complexity scales linearly with the number of predicted objects in the scene. On the contrary, one-stage object detectors are known to be able to perform real-time object detection. In particular, recent papers focus on anchor-free architectures, with the most known probably being the modern versions of YOLO [15, 14, 17].
The first CL approach proposed for a CLOD scenario work is  [18]. The authors propose to use the Learning without Forgetting (LwF) regularization-based approach in the context of a two-stage object detector, Fast-R-CNN [19]. LwF [20], following the idea of knowledge distillation [21], forces the outputs of the training model, called student, to be similar to the corresponding outputs of a frozen version of the model trained on old tasks, called teacher. Following this work, many subsequent approaches in the literature are distillation-based approaches [8].
However, all of them consider either two-stage object detectors [19, 22] or old one-stage object detectors like YOLOv3 [23], which are not anchor-free like FCOS [24] or more recent versions of YOLO like v8 and v10 [17, 14].
More recent works applied to anchor-free fully convolutional detectors [25] revisit regularization-based algorithms for the FCOS object detector. Indeed, this type of model has a separate branch for regression outputs that introduces noise when training with techniques like LwF. Building on this work, a solution for CLOD on the Edge called Latent Distillation was proposed [26]. However, these solutions are tailored FCOS-based architectures, which makes its adoption in modern YOLO architectures challenging.
@@
Also, Replay is used in the CLOD literature[8]. This method stores a portion of the old data in a memory buffer. While training on a new task, the new data is combined with the data from the memory buffer to maintain knowledge of old tasks. According to the CL literature for image classification, Replay is one the most effective solutions for catastrophic forgetting [4, 28].
However, this is not the case in object detection, where task interference arises caused by the missing annotations. When training on a new task, images saved in the memory buffer during previous tasks may also present unlabeled instances of new classes. This means that for those buffer samples, the ground truth interferes with the new class knowledge that the model is trying to learn.
To address these issues, the authors of [12] building on the YOLOv3 architecture proposed to mask the loss associated with new classes during the training of old samples and vice-versa for the new data. However, the masking loss only partially solves the problem caused by the missing annotations. Indeed, this approach fails to fully leverage the potential information contained within the replay memory. Indeed, when saving the current task training samples in the replay memory, the images are labeled only on the classes of the current task but might contain instances of new classes (example in Fig. 3), leading to an under-utilization of the replay memory’s potential.
@@
Figure 2:
3 Replay Consolidation with Label Propagation for Object Detection
To tackle the complexities of Object Detection within a Continual Learning framework, we combine multiple components to maximize their effectiveness and achieve optimal results.
The first component of our method is a Replay memory, but instead of randomly selecting the images for the memory buffer, we adopt the OCDM selection mechanism [10], which aims to have a memory buffer with a more balanced class distribution (see Fig. 4(b)). This selection mechanism is advantageous because it is tailored for settings like CLOD, where each image is associated with multiple labels. The second component is the Label Propagation mechanism, which fully leverages the memory buffer potential by integrating old and new knowledge into its ground truths, thus addressing the interference problems inherent in Replay. To further reduce the task interference problem, we introduce a Masking Loss specifically designed for the anchor-free YOLOv8 model, as explained in Sec. 3.2. Finally, to further decrease the forgetting of old knowledge, we also apply the Feature Distillation technique [11] as presented in Sec. 3.3.
3.1 Label Propagation
In multi-label classification and object detection, replay-based methods commonly face interference issues. This interference arises between the current task samples and the replayed samples when there is an overlap between task classes, as each sample ground truth only has information pertinent to the labels of the associated task. Moreover, the full potential of Replay buffers is not realized in these scenarios, as each sample is stored in the memory buffer with ground truth labels limited to the specific classes of its original task (see Fig. 3(a)).
Label Propagation, illustrated in Fig. 2, is composed of the forward and backward steps.
@@ 3.1 Label Propagation
In the forward step, information from old labels is added to the new data during training and subsequently into the replay buffer through a pseudo-labeling technique. In detail, let $D_{t}=\{X_{t},Y_{t}\}$ be the data of the current task $t$ with labels corresponding to a set of classes $C_{t}$. During the training of task $t$, for a sample $(x,y)\in D_{t}$, the forward step modifies the ground truth $y$ adding pseudo-labels of the classes $C_{1},\ldots,C_{t-1}$ by leveraging the knowledge of the previously trained model $f_{\theta_{t-1}}$. Indeed, the model $f_{\theta_{t-1}}$ provides predictions associated with classes $C_{1},\ldots,C_{t-1}$.
@@
Figure 3: Comparison of the memory for Replay and for RCLP. The vertical axis represents the stored samples of each task, while the horizontal axis represents the objects associated with each class $c_{i}$. (a) Each task of the Replay memory has information only on the classes seen during its iteration. (b) Using the Label Propagation mechanism of RCLP, the saved samples are more informative, containing knowledge of new and old classes.
@@
(b) OCDM Memory
Figure 4: Example showing the difference between Replay and OCDM by comparing the class distribution of the memory buffer. The vertical axis represents the label frequency, while the horizontal axis represents the different labels. (a) On the left, the distribution of Replay is unbalanced. (b) At right, the distribution of OCDM is more balanced because of the selection mechanism.
After training on task $t$, a subset of the samples from task $t$ is stored in the memory buffer $M$. This method ensures that, after training on task $t$, the memory buffer retains samples with information relevant to all tasks up to and including task $t$. On the other hand, the backward step focuses on consolidating the new knowledge gained from training the model on the latest task into the old samples stored in the replay buffer. Specifically, using the new model trained on task $t$, pseudo-labels corresponding to the current task are added to the ground truths of old samples of the memory buffer.
3.2 Masking Loss
Figure 5: Image from replay memory: class "tennis racket" is a new class, while class "person" is an old one. Since the only ground truth available is the one for the "person" class, any model prediction for the tennis racket would be penalized in the classification loss computation.
Label Propagation considerably reduces task interference in replay-based techniques. However, the issue of handling new classes for replay memory samples when training on new tasks remains. On the other hand, since YOLO uses Task Alignment Learning (TAL) [29] for label-prediction assignment, the extent to which interference affects the performance depends on the saved images. In fact, for a given image and its corresponding ground truths, the YOLO assigner assigns at least one model prediction to each ground truth based on the intersection area between the ground truths and the predictions. Consequently, if an object of a new class is not close to an object of an old class, the prediction for the new class object is not assigned to any ground truth, meaning that this prediction doesn’t contribute to the loss.
Therefore, depending on the specific dataset, if instances of new-class objects don’t overlap frequently with those from previous tasks, interference does not arise; otherwise, as in Fig. 5 where, without any class masking, the tennis racket prediction would be penalized in the loss computation, samples from the replay memory can prevent the model from learning new classes effectively.
To tackle this problem independently of the dataset, we propose a masking loss approach similar to the method used in [12]. In this method, the contribution of new classes for computing the classification loss is ignored for the samples saved in the memory buffer; recalling that, for a prediction assigned to a ground truth, the classification loss is given by $\mathcal{L}_{\text{cls}}=\sum_{i}\,\mathcal{L}_{\text{cls}}^{i}$, where $\mathcal{L}_{\text{cls}}^{i}$ is the usual Binary Cross Entropy for class $i$, by applying the masking approach, the classification loss for any sample in the Replay memory is given by:
3.3 Feature Distillation
Finally, to further enhance the method’s performance, we employ the Feature Distillation technique [11], similar to RCLP and SID approaches[25]. The core idea behind feature distillation is that the features generated by the new model for both new and old data should remain close to those produced by the previous model for the same data. This approach helps the model maintain a more cohesive memory, mitigating the effects of catastrophic forgetting and ensuring more stable performance across tasks.
In our method, we perform feature distillation on the model’s intermediate representations. In details, let our model $f(x)=g_{\phi}(h_{\omega}(x))$, where given an input $x$, $h$ gives the intermediate features, and $g$ produces the model output based on such features. Formally, given an input sample $x$, the distillation loss on intermediate features is given by:

### v2

@@ Abstract
Continual Learning (CL) aims to learn new data while remembering previously acquired knowledge. In contrast to CL for image classification, CL for Object Detection faces additional challenges such as the missing annotations problem. In this scenario, images from previous tasks may contain instances of unknown classes that could reappear as labeled in future tasks, leading to task interference in replay-based approaches. Consequently, most approaches in the literature have focused on distillation-based techniques, which are effective when there is a significant class overlap between tasks. In our work, we propose an alternative to distillation-based approaches with a novel approach called Replay Consolidation with Label Propagation for Object Detection (RCLPOD). RCLPOD enhances the replay memory by improving the quality of the stored samples through a technique that promotes class balance while also improving the quality of the ground truth associated with these samples through a technique called label propagation. RCLPOD outperforms existing techniques on well-established benchmarks such as VOC and COC. Moreover, our approach is developed to work with modern architectures like YOLOv8, making it suitable for dynamic, real-world applications such as autonomous driving and robotics, where continuous learning and resource efficiency are essential.
Keywords Continual Learning  $\cdot$ Object Detection  $\cdot$ Deep Learning
@@ 1 Introduction
Object Detection is a critical computer vision problem involving both the localization and classification of objects within an image. However, a fundamental limitation of Deep Learning systems is their inability to learn incrementally over time, as they are prone to Catastrophic Forgetting [1, 2] when exposed to new data. Continual Learning (CL) is a research field focused on enabling models to learn new information while retaining previously acquired knowledge. Applications such as robotics [3] and autonomous driving [4] can benefit greatly from models capable of adapting and learning continuously from streaming data.
Most CL studies for image classification, have shown that replay-based approaches are the most effective method to achieve strong results [5, 6, 7, 8]. However, in the Continual Learning for Object Detection (CLOD) setting, replay-based approaches suffer from the missing annotations problem [9]. For example, the image in Fig. 1 contains the classes toy, mug, and book, but not all of these objects are consistently labeled across tasks. This causes task interference and memory efficiency issues, making replay-based approaches underperform.
To address the limitations of replay approaches for CLOD, we propose a novel technique called Replay Consolidation with Label Propagation for Object Detection (RCLPOD). RCLPOD achieves superior performance on well-known benchmarks like COCO and VOC, demonstrating that replay-based approaches can perform as well or even better than distillation-based methods. RCLPOD improves the replay memory promoting class balance while also improving the quality of the ground truth of the stored samples through a technique called Label Propagation (see Fig. 2 for an overview of the RCLPOD framework).
Another limitation in the current CLOD landscape is that most studies are based on models like Faster R-CNN [10, 11] or YOLO architectures such as YOLOv3 [12] and v5 [13]. However, these models differ significantly in terms of architecture design from recent YOLO versions. This paper provides a benchmark for modern YOLO architectures by testing several CL methods using the anchor-free YOLOv8 architecture [14]. Therefore, this work can serve as a valuable benchmark for future research on modern YOLO architectures within the CLOD framework.
Figure 1: Continual Learning for Object Detection pipeline. The model learns to detect new classes at each incremental training stage (task). However, the scenario is challenging due to the missing annotations problem. The example in the current task might have been learned as background in previous tasks (e.g. the class toy is shown in Task 1 but not in Task 2 or 3.).
@@
  • We propose a novel approach called RCLPOD, which addresses key limitations of replay techniques in the CLOD setting.
  • We conduct extensive experiments with RCLPOD, demonstrating its superior performance on the established VOC and COCO CLOD benchmarks.
  • We benchmark our approach and other CL techniques using more modern versions of YOLO than those commonly evaluated in the literature.
The remainder of the paper is structured as follows. Sec. 2 provides an overview of the CL literature focusing on CLOD. In Sec. 3, we describe our novel approach, RCLPOD. Sec. 4 describes the experimental setting, including the considered metrics, evaluated scenarios, and experimental details. In Sec. 5, we show and discuss the results. Finally, Sec. 6 concludes this work by discussing limitations and future research directions.
@@ 2 Related Work
Continual Learning aims to enable DL models to continuously accumulate knowledge over time without the need to re-train from scratch and with limited access to past data [15]. In particular, this framework has emerged to tackle the Catastrophic Forgetting problem [1, 2]; after a DL model is trained on one task and subsequently trained on another, it tends to forget the previously learned task.
Most of the literature on Continual Learning for Object Detection focuses on the Class Incremental Learning (CIL) scenario, where each new task introduces new classes [9]. For example, Fig. 1 depicts a CLOD setting composed of three tasks where a new class must be learned each time. At the end of the training, the model should be able to provide predictions on all these classes. Most of the works in CLOD literature were conducted on two-stage object detectors, whose time complexity scales linearly with the number of predicted objects in the scene. On the contrary, one-stage object detectors are known to be able to perform real-time object detection. In particular, recent papers focus on anchor-free architectures, with the most known probably being the modern versions of YOLO [16, 14, 17].
The first CL approach proposed for a CLOD scenario was proposed in [18]. The authors propose to use the Learning without Forgetting (LwF), a regularization-based approach, in the context of a two-stage object detector, Fast-R-CNN [19]. LwF [20], following the idea of knowledge distillation [21], forces the outputs of the training model, called student, to be similar to the corresponding outputs of a frozen version of the model trained on old tasks, called teacher. Following this work, many subsequent in the literature are distillation-based approaches [9].
However, all of them consider either two-stage object detectors [19, 22] or one-stage object detectors like YOLOv3 [23], which are not anchor-free like fully convolutional one-stage object detector (FCOS) [24] or more recent versions of YOLO like v8 [14].
More recent works applied to anchor-free fully convolutional detectors [25] revisit regularization-based algorithms for the FCOS object detector. In particular, they address the noise introduced by the separate branch for regression outputs. Building on this work, a solution for CLOD on the Edge called Latent Distillation was proposed [26]. However, these solutions are tailored FCOS-based architectures, which makes its adoption in modern YOLO architectures challenging.
@@
The CLOD literature includes also Replay [9]. This method stores a portion of the old data in a memory buffer. While training on a new task, the new data is combined with the data from the memory buffer to maintain knowledge of old tasks [28]. According to the CL literature for image classification, Replay is one the most effective solutions for catastrophic forgetting [5, 29].
However, this is not the case in object detection, where task interference arises caused by the missing annotations.
To address these issues, the authors of [12] building on the YOLOv3 architecture proposed to mask the loss associated with new classes during the training of old samples and vice-versa for the new data. However, the masking loss only partially solves the problem caused by the missing annotations. This approach fails to fully leverage the potential information contained within the replay memory (memory efficiency issue).
Contrary to the previously cited works, anchor-free one-stage object detectors are the most used in practical applications nowadays. Therefore, we propose to benchmark our method using the widely adopted YOLOv8, showing how our approach mitigates the problems that affect either distillation-based solutions or replay ones.
3 Replay Consolidation with Label Propagation for Object Detection
In this work, a novel approach is introduced called Replay Consolidation with Label Propagation for Object Detection (RCLPOD), aimed at improving Continual Learning in Object Detection by addressing task interference and enhancing memory replay.
@@ 3 Replay Consolidation with Label Propagation for Object Detection
Figure 2: Scheme of the RCLPOD method: (a) During the training of the new task, the new samples are first processed through the old model to generate pseudo-labels. Then they are combined with the samples from the replay memory to update the model, considering also additional losses like the Masking Loss and the one associated with feature distillation to reduce the drift of old intermediate representations. (b) Post-training procedure to update the replay memory. This procedure enhances the replay memory via label propagation. The backward step updates the old samples with new knowledge, while the forward step integrates old knowledge into new task samples. Some of the new enriched samples are stored in the memory buffer.
One of the main issues of the CLOD setting is the missing annotations problem. In the CLOD setting, images seen in previous tasks may contain unlabeled classes that the model must learn as new classes in future tasks, causing missing annotations [9]. For example, the image in Fig. 1 has the classes person, dog, and car. However, this image has only one labeled class for each task. Then, this image could be saved multiple times in the replay memory, each time with different annotations based on the task.
This problem impacts negatively on the performance of Replay because it causes the task interference and the memory efficiency issues. The task interference arises between the current task samples and the replayed samples thar contain unlabeled instances of new classes. This means that for those buffer samples, the ground truth interferes with the new class knowledge that the model is trying to learn
Moreover, even if such a problem was solved (e.g. through a Masking Loss [12]), such an approach would still have the memory efficiency issue. This is because the full potential of the replay buffer is not realized, as each sample is stored in the memory buffer with ground truth labels limited to the specific classes of its original task (see Fig. 3(a)).
3.1 Replay Memory with Label Propagation
To solve the discussed issues of the replay-based approaches in the CLOD setting, we propose to implement a technique never considered before for CLOD called Label Propagation (LP) mechanism. Originally proposed for the multi-label image classification setting, LP aims to enrich the replay memory by adding pseudo-labels to old and new samples [30] (see Fig. 3). Adding pseudo-labels of previously learned classes to new task samples enriches the replay memory, allowing the model to retain knowledge of old classes while learning new ones.
While multi-label image classification and object detection differ significantly in terms of model architecture and training process, we argue that LP can be exploited to enhance the performance of CLOD.
@@ 3.1 Replay Memory with Label Propagation
We implement LP for CLOD considering the procedure structure into two different steps (see Fig. 2(b)). In the forward step, information from old labels is added to the new data during training and subsequently into the replay buffer through a pseudo-labeling technique. In detail, let $D_{t}=\{X_{t},Y_{t}\}$ be the data of the current task $t$ with associated a set of classes $C_{t}$ that appears in task $t$. During the training of task $t$, for a sample $(x,y)\in D_{t}$, the forward step modifies the ground truth $y$ adding pseudo-labels of the classes $C_{1},\ldots,C_{t-1}$ by leveraging the knowledge of the previously trained model $f_{\theta_{t-1}}$. Indeed, the model $f_{\theta_{t-1}}$ was trained with classes $C_{1},\ldots,C_{t-1}$. This method ensures that, after training on task $t$, the memory buffer retains samples with information relevant to all tasks up to and including task $t$.
On the other hand, the backward step focuses on consolidating the new knowledge gained from training the model on the latest task into the old samples stored in the replay buffer. Specifically, using the new model trained on task $t$, pseudo-labels corresponding to the current task are added to the ground truth of old samples in the memory buffer.
3.2 Balanced Replay Memory
To improve the performance of RCLPOD, we propose to deviate from the classic selection mechanism adopted by Replay. However, such a decision is sub-optimal since object detection datasets are normally unbalanced, like the case for VOC and COCO datasets. Therefore, we propose enhancing the quality of the samples in the replay memory by ensuring a more balanced representation of each class.
Specifically, we propose to use a selection mechanism called OCDM [31], which aims to have a memory buffer with a more balanced class distribution (see Fig. 4(b)). This selection mechanism is advantageous because it is tailored for settings like CLOD, where each image is associated with multiple objects. Giving higher priority to images with underrepresented classes helps achieve more balance and promotes class fairness. Specifically, a greedy approach is applied which removes samples one by one based on how much they contribute to the target uniform distribution (see Sec. A.7 for the pseudocode).
In the CLOD context, OCDM and Label Propagation work in synergy to optimize the replay memory, as demonstrated by the conducted ablation study of Sec. 5.3.
If OCDM is applied without Label Propagation, it may need to select the same samples multiple times to achieve balanced class representation, which could limit the diversity of the samples in the buffer. However, with Label Propagation enhancing each sample with pseudo-labels for additional classes, the samples in memory become more informative. This allows OCDM to select a more varied set of samples, improving the overall representativeness of the replay memory while reducing redundancy.
3.3 Masking overlapping objects
Label Propagation considerably reduces task interference in replay-based techniques. However, the issue of handling new classes for replay memory samples when training on new tasks remains. On the other hand, since YOLO uses Task Alignment Learning (TAL) [32] for label-prediction assignment, the extent to which interference affects the performance depends on the saved images. In fact, YOLO’s assigner links each ground truth with at least one model prediction based on the intersection area. Thus, if a new-class object is distant from old-class objects, no ground truth is assigned to its prediction, and it doesn’t affect the loss. Consequently, when new-class objects rarely overlap with old ones, interference is minimized. When instead the dataset presents frequent overlaps between objects, as in Fig. 5, interference arises. We define this as the overlapping objects issue.
To tackle this problem independently of the dataset, inspired by [12], we propose a custom masking loss approach for YOLOv8. In particular, we propose to ignore the contribution of new classes for the classification loss computation for every prediction associated to any sample saved in the memory buffer; recalling that, for a prediction assigned to a ground truth, the classification loss is given by $\mathcal{L}_{\text{cls}}=\sum_{i}\,\mathcal{L}_{\text{cls}}^{i}$, where $\mathcal{L}_{\text{cls}}^{i}$ is the usual Binary Cross Entropy for class $i$, the classification loss for any sample in the Replay memory is given by:
@@ 3.3 Masking overlapping objects
Figure 3: Comparison of the memory storage for Replay and RCLPOD. The vertical axis indicates the stored samples of each task, while the horizontal axis represents the objects associated with each class $c_{i}$. (a) Each task of the Replay memory has information only on the classes seen during its iteration. (b) Using the Label Propagation mechanism, the saved samples are more informative, containing knowledge of new and old classes.
@@
(b) RCLPOD Memory
Figure 4: An example comparing the class distributions in the memory buffer of Replay and RCLPOD. The vertical axis indicates the label frequency, while the horizontal axis represents the different labels. (a) On the left, Replay shows an unbalanced distribution. (b) At right, the distribution of RCLPOD is more balanced because of the selection mechanism.
Figure 5: Example of task interference for YOLO architecture due to overlapping objects. Image from replay memory: class "tennis racket" is a new class, while class "person" is an old one. Since the only ground truth available is the one for the "person" class, any model prediction for the tennis racket would be penalized in the classification loss computation.
3.4 Improving Stability
By employing the RCLPOD method as described above, we are able to provide more information to the replayed samples improving both stability and plasticity when learning a new CL task. To further enhance model stability, we aim to minimize the drift of intermediate representations of previously learned objects. However, to balance the stability-plasticity trade-off, we consider low-level features, as they are more generalizable and can be shared across tasks. Therefore, to improve the stability of the model and reduce forgetting, we employ a Feature Distillation (FD) technique [33] by aligning the low-level features of the new and old models.
Specifically, for a YOLOv8 architecture, we perform distillation on the model’s intermediate representations as follows: Let our model $f(x)=g_{\phi}(h_{\omega}(x))$, where for an input $x$, $h$ gives the intermediate features, and $g$ produces the model output based on such features. Formally, given an input sample $x$, the distillation loss on intermediate features is the following: